# Text2Cypher: Data Pruning using Hard Example Selection

### v1

@@ Abstract.
Accepted to LLM-DPM 2025 workshop (in conjunction with SIGMOD/PODS 2025). This is the submitted version; a revised version incorporating reviewer feedback will be shared later.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Query languages for non-relational enginesccs: Information systems Graph-based database modelsccs: Information systems Data cleaningccs: Computing methodologies Machine translationccs: Computing methodologies Supervised learning
@@ 3.1. Selecting Challenging Instances
  • Complexity-Based Hard-Example Selection: Previous analysis of Neo4j (Ozsoy, 2025) identified data sources and databases where fine-tuned models struggled most. Based on this analysis: (i) The chosen databases are three demonstration databases of Neo4j  [^1] Neo4j Text2Cypher Crowdsourcing App: https://text2cypher.vercel.app/  [^2] Neo4j Browser Demo: https://demo.neo4jlabs.com:7473/browser/, namely ”recommendations, companies, neoflix”, and (ii) The selected data-sources are: ”functional_cypher”, ”synthetic_gemini”, and ”text2cypher2023_train”.
@@ 4.1. Experimental Setup and Evaluation Metrics
We fine-tuned a baseline model, ’unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit’, using various training datasets prepared according to the proposed hard-example selection methods. During evaluation, we used the test set and fine-tuned models to generate Cypher queries based on input natural language questions and corresponding database schemas. After generating the Cypher queries, we applied a post-processing step to remove unwanted text, such as the ’cypher:’ prefix. Details of the fine-tuning setup and parameters are provided in Appendix A.
@@ 5. Conclusion
Appendix A Fine-tuning Parameters
For fine-tuning, we used a RunPod (RunPod, 2024) GPU environment with a single A40 machine. The fine-tuning process was conducted using the Unsloth(Unsloth, 2024) framework. The following parameters were used for fine-tuning are presented in Table 2.
@@ Appendix A Fine-tuning Parameters
$load\_in\_4bit:True$,
$truncation\_side:"left"$,
$padding\_side:"left"$
@@
$lora\_alpha:16$,
$lora\_dropout:0$,
$bias:"none"$,
$random\_state:3407$,
$use\_rslora:False$,
$loftq\_config:None$
@@
$per\_device\_train\_batch\_size:2$,
$gradient\_accumulation\_steps:4$,
$warmup\_steps:5$,
$num\_train\_epochs:1$,
$learning\_rate:2e-4$,
$fp16:notis\_bfloat16\_supported()$,
$bf16:is\_bfloat16\_supported()$,
$optim:"adamw\_8bit"$,
$weight\_decay:0.01$,
$lr\_scheduler\_type:"linear"$,

### v2

@@ Abstract.
Accepted to LLM-DPM 2025 workshop (in conjunction with SIGMOD/PODS 2025)
copyright: acmlicensedjournalyear: 2025conference: LLM-DPM 2025 workshop (in conjunction with SIGMOD/PODS 2025); 2025; Berlin, Germanyccs: Information systems Query languages for non-relational enginesccs: Information systems Graph-based database modelsccs: Information systems Data cleaningccs: Computing methodologies Machine translationccs: Computing methodologies Supervised learning
@@ 3.1. Selecting Challenging Instances
In our previous work (Ozsoy, 2025), we have executed a comprehensive analysis of model performance on the Neo4j Text2Cypher (2024) dataset (Ozsoy et al., 2025). This analysis explored evaluation results from multiple perspectives, including key metrics (such as Google-Bleu and Exact Match), assigned complexity levels, and breakdowns by data source, database type, and fine-tuned model. The statistical analyses (e.g., averages, standard deviations) revealed consistent patterns of model struggle, particularly on examples from specific databases and data sources. Further error analysis highlighted that these challenges were often due to inconsistencies in ground-truth Cypher queries (such as varying use of WHERE clauses or aggregation methods), limitations of existing evaluation metrics, and underlying model weaknesses. These findings directly motivated our hard-example selection strategies, which aim to construct a more informative and targeted training subset by focusing on the most challenging instances.
@@
  • Complexity-Based Hard-Example Selection: Guided by our analysis (Ozsoy, 2025), we identified data sources and databases where fine-tuned models struggled most. Based on this analysis: (i) The chosen databases are three demonstration databases of Neo4j  [^1] Neo4j Text2Cypher Crowdsourcing App: https://text2cypher.vercel.app/  [^2] Neo4j Browser Demo: https://demo.neo4jlabs.com:7473/browser/, namely ”recommendations, companies, neoflix”, and (ii) The selected data-sources are: ”functional_cypher”, ”synthetic_gemini”, and ”text2cypher2023_train”.
@@ 4.1. Experimental Setup and Evaluation Metrics
We fine-tuned a baseline model, ’unsloth/Meta-Llama-3.1-8B-Instruct-bnb-4bit’, using various training datasets prepared according to the proposed hard-example selection methods. During evaluation, we used the test set and fine-tuned models to generate Cypher queries based on input natural language questions and corresponding database schemas. After generating the Cypher queries, we applied a post-processing step to remove unwanted text, such as the ’cypher:’ prefix. Details of the fine-tuning setup and parameters are provided in Appendix B.
@@ 5. Conclusion
Appendix A Declaration on Generative AI Usage
During the preparation of this work, the author(s) used Chat-GPT in order to: ’Improve writing style’ and ’Paraphrase and reword’. After using these tool(s)/service(s), the author(s) reviewed and edited the content as needed and take(s) full responsibility for the publication’s content.
Appendix B Fine-tuning Parameters
For fine-tuning, we used a RunPod (RunPod, 2024) GPU environment with a single A40 machine. The fine-tuning process was conducted using the Unsloth(Unsloth, 2024) framework. The parameters used for fine-tuning are presented in Table 2.
@@ Appendix B Fine-tuning Parameters
load_in_4bit:Trueload\_in\_4bit:True,
truncation_side:lefttruncation\_side:"left",
padding_side:leftpadding\_side:"left"
@@
lora_alpha:16lora\_alpha:16,
lora_dropout:0lora\_dropout:0,
bias:nonebias:"none",
random_state:3407random\_state:3407,
use_rslora:Falseuse\_rslora:False,
loftq_config:Noneloftq\_config:None
@@
per_device_train_batch_size:2per\_device\_train\_batch\_size:2,
gradient_accumulation_steps:4gradient\_accumulation\_steps:4,
warmup_steps:5warmup\_steps:5,
num_train_epochs:1num\_train\_epochs:1,
learning_rate:2e4learning\_rate:2e-4,
fp16:notis_bfloat16_supported()fp16:notis\_bfloat16\_supported(),
bf16:is_bfloat16_supported()bf16:is\_bfloat16\_supported(),
optim:adamw_8bitoptim:"adamw\_8bit",
weight_decay:0.01weight\_decay:0.01,
lr_scheduler_type:linearlr\_scheduler\_type:"linear",